OpenStudy (anonymous):

two bicyclists, Jack and Jill, are taking the same route. Jack leaves at 7:00 a.m. at a constant speed of 10 mph. Jill leaves at 7:30 a.m. also at a constant speed. At 7:45 a.m., Jack is 5.5 miles ahead of Jill. a) How fast is Jill going? b) How long will it take her to travel 40 miles

OpenStudy (anonymous):

(a) 8 miles per hour (b) 5 hours

OpenStudy (anonymous):

can you show me how did you get the answers ahan98 ?

OpenStudy (anonymous):

7:00 to 7:45 = 45 minutes 45 minutes = 3/4 of an hour Jack traveled 7.5 miles (3/4 * 10) Then 7:30 - 7:45 = 15 minutes (1/4 of an hour) Jill traveled (7.5 - 5.5 = 2) miles. 2 * 4 = 8 mph For (b) You have distance = rate * time 40 = 8 * time time = 5 hours

OpenStudy (anonymous):

Got it yet?